OpenStudy (anonymous):

evaluate y^ii(2),when y=x^-2 a 0.375 b 0.148 c 0.268 d 0.425

OpenStudy (jamesj):

If y = 1/x^2, what is y' and y'' in general? That's the first step.

OpenStudy (anonymous):

0/0 but i see y^ii (2) how about that?

OpenStudy (jamesj):

Still there? 0/0 has nothing to do with it. The first derivative is \[ y' = -2.x^{-3} \] What's the second derivative?

OpenStudy (anonymous):

6x

OpenStudy (jamesj):

No, \[ y'' = (-3)(-2)x^{-4} = 6x^{-4} \] Hence you can now evaluate \[ y''(2) \] by substituting \( x = 2 \) into the formula for the second derivative.

OpenStudy (anonymous):

wow thanks the answer is near letter d is this correct?

OpenStudy (jamesj):

No. Substitute x=2 into that equation and you have \[ y''(2) = \frac{6}{16} = \frac{3}{8} \] Which decimal answer is that equal to?

OpenStudy (jamesj):

Are you following?

OpenStudy (anonymous):

0.375

OpenStudy (anonymous):

sorry i thought what i did is my prof do.he times 6 in the sqrt because for him x=1